A progressive wave travelling along a taut string is given by

Y= 25mm sin 2đťś‹ (t/20ms+x/40m)

i. Find the phase velocity of the wave.

ii. Give an equation of a standing wave that can be formed from the 

above progressive wave.


1
Expert's answer
2021-03-17T16:55:34-0400

(a) The general equation of the progressive wave can be written as follows:


"y(x,t)=Asin2\pi(\dfrac{x}{\lambda}+\dfrac{t}{T}),"

here, "\lambda" is the wavelength of the wave, "T" is the period of the wave.

As we can see from the equation of the progressive wave, "\lambda=40\ m, T=20\ s."

Then, we can find the phase velocity of the wave as follows:


"v_p=\dfrac{\lambda}{T}=\dfrac{40\ m}{20\ s}=2\ \dfrac{m}{s}."

(b) Let's rewrite our equation of progressive wave:


"y(x,t)=25sin2\pi(\dfrac{x}{40}+\dfrac{t}{20}),""y(x,t)=25sin(0.05\pi x+0.1\pi t)."

Let's consider two identical waves that move in opposite directions. The first wave has a wave function of "y_1(x,t)=25sin(0.05\pi x+0.1\pi t)" and the second wave has a wave function "y_2(x,t)=25sin(0.05\pi x-0.1\pi t)". The waves interfere and form a resultant wave:


"y(x,t)=y_1(x,t)+y_2(x,t),""y(x,t)=25sin(0.05\pi x+0.1\pi t)+25sin(0.05\pi x-0.1\pi t)."

Using the trigonometric identity


"sin(\alpha \pm \beta)=sin\alpha cos\beta \pm cos\alpha sin\beta"

we can write the equation representing the standing wave:


"y(x,t)=2\cdot25sin(0.05\pi x)cos(0.1\pi t),""y(x,t)=50sin(0.05\pi x)cos(0.1\pi t)."